how to find the mean of x ie E(x)


1
Expert's answer
2021-12-20T16:54:35-0500

The mean of x can be found by the formula: "\frac{\left(\sum \:x_i\right)}{n}" where "n" is the number of terms and "\sum\:\:x_i" is the sum of the terms.